A rectangular floor has a perimeter of 54 feet. If the length is 3 more than the width, what are the dimensions of the floor?
1
Expert's answer
2012-11-29T08:02:15-0500
P=54=2(w+l)
Let x feet - width then x+3 - length
P=2(x+3+x)=4x+6=54
x=12 feet
Width 12 feet, length 15 feet